### v1

@@ Abstract
With the development of smart cities, the demand for continuous pedestrian navigation in large-scale urban environments has significantly increased. While global navigation satellite systems (GNSS) provide low-cost and reliable positioning services, they are often hindered in complex urban canyon environments. Thus, exploring opportunistic signals for positioning in urban areas has become a key solution. Augmented reality (AR) allows pedestrians to acquire real-time visual information. Accordingly, we propose a low-cost visual-inertial positioning solution. This method comprises a lightweight multi-scale group convolution (MSGC)-based visual place recognition (VPR) neural network, a pedestrian dead reckoning (PDR) algorithm, and a visual/inertial fusion approach based on a Kalman filter with gross error suppression. The VPR serves as a conditional observation to the Kalman filter, effectively correcting the errors accumulated through the PDR method. This enables the entire algorithm to ensure the reliability of long-term positioning in GNSS-denied areas. Extensive experimental results demonstrate that our method maintains stable positioning during large-scale movements. Compared to the lightweight MobileNetV3-based VPR method, our proposed VPR solution improves Recall@1 by at least 3% on two public datasets while reducing the number of parameters by 63.37%. It also achieves performance that is comparable to the VGG16-based method. The VPR-PDR algorithm improves localization accuracy by more than 40% compared to the original PDR.
@@
A lightweight multi-scale group convolutional network (MSGC-Net) for visual beacon recognition
@@
A novel visual/inertial integrated localization scheme with gross error suppression method
Detailed performance analysis for the MSGC-Net
Detailed performance analysis for the PDR method
@@ 1 Introduction
VPR is typically approached as an image retrieval task Arandjelovic et al. (2016); Hausler et al. (2021); Yu et al. (2019); Xu et al. (2021). Given a query image of the target position, the system retrieves the best-matching reference image with a pre-installed position by traversing a database. With the establishment and enrichment of urban geographic databases, such as street view maps, VPR has become a powerful opportunistic signal for urban localization. The central challenge is to create a compact image descriptor to represent an image effectively. Among all the methods, the aggregation-based techniques have demonstrated superior performance. Inspired by traditional VLAD Jégou et al. (2010), NetVLAD Arandjelovic et al. (2016) has emerged as the most prominent global image representation for VPR. In pursuit of even better performance, researchers have enhanced NetVLAD with additional functional modules Peng et al. (2021a); Yu et al. (2019); Peng et al. (2021b). Although these variants achieve state-of-the-art results, they rely on large backbone networks such as VGG16 Simonyan and Zisserman (2014) for feature extraction. The complexity of these deep architectures increases both memory usage and computational costs, placing greater demands on hardware devices. In addition, the VPR algorithm only provides discrete position estimation instead of a continuous trajectory.
@@
  • We design a lightweight VPR neural network, MSGC-NetVLAD, that provides absolute position references for the PDR algorithm.
  • To ensure reliable long-term positioning, we integrate the PDR and VPR systems using a Kalman filter which offers trajectory smoothing and gross error suppression.
@@ 2.3.4 Position Propagation
where $\mathbf{p}_{i}=\left[x_{i},y_{i}\right]^{\text{T}}$ is the 2D position at $i$-th step; $L_{i}$ is the step length computed by equation (11) at $i$-th step; $\psi_{i}$ is the heading computed by equation (14) at $i$-th step.
@@ 3.1 Dataset and Evaluation Metric
For the final evaluation of the proposed visual-inertial localization algorithm, we incorporated the Real-World Walk dataset, which includes two trajectories along with time-aligned IMU data, discrete images, and ground truth positions.
Consistent with standard evaluation protocols for these datasets, we measured VPR model performance using Recall@N Arandjelovic et al. (2016). Additionally, the efficiency of the proposed VPR model was assessed through model parameter quantity and floating-point operations per second (FLOPS). To evaluate the efficacy of our visual-inertial localization algorithm, GPS benchmarks were employed.
@@ 3.4 PDR-VPR Localization Performance Evaluation
  • Trajectory 2, Dense Feature Environment: This was conducted in a typical high-feature environment, where the participant walked a closed loop among densely packed buildings. The environment offered numerous available visual features for positioning.
We begin by presenting the evaluation results obtained in the lakeside environment, as illustrated in Figure 8. These results include both the proposed PDR trajectory and the fused PDR-VPR trajectory. In the sparse feature environment, we set a total of 11 visual beacons, of which 7 were successfully recognized (marked as Red Points) while 4 were unrecognized (marked as Green Points).
We aim to clarify the factors influencing this result. As shown in Figure 4, we provide a comparison of the actual landmark images corresponding to four unrecognized points. The unrecognized points may be attributed to several factors: the walking route actually passed through the visual beacons (Points 3 and 4), as well as a lack of distinctive features (Points 1 and 2). Additionally, to demonstrate the effectiveness and necessity of our gross error suppression threshold, we separately present the trajectory results obtained by directly selecting the image with the highest recall@1 score as the correct match. As shown in the Figure 5, without PDR prior information as a constraint, using image similarity alone as the recognition criterion leads to completely unusable localization results.
This is due to the limited Recall@1 accuracy of the lightweight VPR network, particularly when the training and testing datasets are from different sources (our method was trained solely on the Pitts30k-train dataset and was not further trained during subsequent testing). The results in Table 3 further support this observation. Once a recognition error occurs, its impact on the subsequent localization results is catastrophic.
With the aid of visual beacons, the fused trajectory aligns more closely with the ground truth compared to the pure PDR trajectory. This improvement is particularly evident in Figure 6, which displays the cumulative distribution function (CDF) curves for the fused trajectory, fused trajectory without gross error suppression threshold, the pure PDR trajectory and the PDR trajectory without MDR, which highlighting the enhancements made.
@@
Figure 5: Trajectory 1 without Gross Error Suppression Threshold
The algorithm demonstrated improved performance in environments with dense features. As illustrated in Figure 9, out of 12 visual beacons, 11 were successfully recognized (indicated by the Red Points), while only one beacon failed to be recognized (shown as the Green Point), resulting in a recognition accuracy of 91.7%. The failure point is shown in Figure 9, where the problematic beacon is situated among numerous repetitive trees, which obscure the nearby architectural features. Denser features result in more successfully recognized beacons, the presence of more successfully recognized visual beacons led to significant enhancements in the fused trajectory. Figure 7 presents the CDF curves of the trajectories within the dense feature environment, showcasing these improvements.
Figure 6: CDF Comparison on Sparse Feature Environment
Figure 7: CDF Comparison on Dense Feature Environment
Figure 8: Trajectory comparison of Trajectory 1 in a sparse feature environment. The red points are correctly recognized, while the green ones are unrecognized. The indices of the green points from bottom to top are 1, 3, and 4.
Figure 9: Trajectory comparison of Trajectory 2 in dense feature environment. The red points are correctly recognized, while the green ones are unrecognized. The index of the green point is 1.
@@ 4 Conclusion
This paper aims to provide stable and continuous positioning in large-scale urban environments by combining opportunistic visual landmark recognition and inertial pedestrian navigation, targeting potential applications on future mobile devices such as AR. To achieve this, we propose a lightweight MSGC-NetVLAD network for visual beacon recognition, which utilizes multi-scale convolutions and group convolutions to ensure a lightweight yet high-accuracy VPR. Additionally, a PDR method based on magnetic disturbance rejection is introduced to ensure positioning continuity. To address gross error issues in VPR, we propose using the natural divergence of the PDR system as a confidence threshold to constrain VPR observation updates. All modules are integrated using a Kalman filter. Extensive experiments demonstrate that our proposed MSGC-NetVLAD network achieves competitive accuracy with only approximately 1.11M parameters, and our overall PDR-VPR fusion algorithm improves the positioning accuracy of traditional PDR by up to 46.86% at a confidence level of 75%. However, the current work is limited to the algorithmic level, and we hope to integrate the proposed algorithm into mobile devices in the future. We believe this research will make a significant contribution to the application of AR technology and the development of smart cities.
@@ Acknowledgements
The authors would like to express their sincere gratitude to the editor and reviewers for their helpful comments throughout the review process, which significantly enhanced the quality of this paper. We also extend our thanks to Ms. Zhou Fengyi for her contributions to the design of the illustration graphs included in this work.

### v2

@@ Abstract
With the development of smart cities, the demand for continuous pedestrian navigation in large-scale urban environments has significantly increased. While global navigation satellite systems (GNSS) provide low-cost and reliable positioning services, they are often hindered in complex urban canyon environments. Thus, exploring opportunistic signals for positioning in urban areas has become a key solution. Augmented reality allows pedestrians to acquire real-time visual information. Accordingly, we propose a low-cost visual-inertial positioning solution. This method comprises a lightweight multi-scale group convolution (MSGC)-based visual place recognition (VPR) neural network, a magnetic disturbance rejection (MDR)-based pedestrian dead reckoning (PDR) algorithm, and a visual/inertial fusion approach based on a Kalman filter with gross error suppression. The VPR serves as a conditional observation to the Kalman filter, effectively correcting the errors accumulated through the MDR-PDR method. This enables the entire algorithm to ensure the reliability of long-term positioning in GNSS-denied areas. Extensive experimental results demonstrate that our method maintains stable positioning during large-scale movements. Compared to the lightweight MobileNetV3-based VPR method, our proposed VPR solution improves Recall@1 by at least 3% on two public datasets while reducing the number of parameters by 63.37%. It even achieves performance that is comparable to the larger VGG16-based method on the Pitts30k-test dataset, all with approximately 1.11 M parameters. In addition, the PDR-VPR algorithm improves 75% localization accuracy by more than 40% compared to the MDR-PDR.
@@
A lightweight multi-scale group convolutional network (MSGC-NetVLAD) for visual place recognition
@@
A novel visual/inertial integrated localization scheme with gross error suppression
Detailed performance analysis for the MSGC-NetVLAD
@@ 1 Introduction
VPR is typically approached as an image retrieval task Arandjelovic et al. (2016); Hausler et al. (2021); Yu et al. (2019); Xu et al. (2021). Given a query image of the target position, the system retrieves the best-matching reference image with a pre-installed position by traversing a database. With the establishment and enrichment of urban geographic databases, such as street view maps, VPR has become a powerful tool to capture opportunistic visual signals for urban localization. These signals refer to the use of occasional, context-dependent visual information—such as landmarks, street views, or fixed visual beacons—to enhance localization systems in environments where traditional GNSS signals are weak or unavailable, such as urban canyons or indoor spaces. These visual signals serve as an alternative positioning method that complements GNSS, allowing for reliable localization even in challenging conditions. The central challenge for VPR is to create a compact image descriptor to represent an image effectively. Among all the methods, the aggregation-based techniques have demonstrated superior performance. Inspired by traditional VLAD Jégou et al. (2010), NetVLAD Arandjelovic et al. (2016) has emerged as the most prominent global image representation for VPR. In pursuit of even better performance, researchers have enhanced NetVLAD with additional functional modules Peng et al. (2021a); Yu et al. (2019); Peng et al. (2021b). Although these variants achieve state-of-the-art results, they rely on large backbone networks such as VGG16 Simonyan and Zisserman (2014) for feature extraction. The complexity of these deep architectures increases both memory usage and computational costs, placing greater demands on hardware devices. In addition, the VPR algorithm only provides discrete position estimation instead of a continuous trajectory.
@@
  • We design a lightweight VPR neural network, MSGC-NetVLAD, that provides absolute position references for the MDR-PDR algorithm.
  • To ensure reliable long-term positioning, we integrate the MDR-PDR and VPR systems using a Kalman filter which offers trajectory smoothing and gross error suppression (GES).
@@ 2.3.4 Position Propagation
where $\mathbf{p}_{i}=\left[x_{i},y_{i}\right]^{\text{T}}$ is the 2D position at $i$-th step; $L_{i}$ is the step length computed by equation (11) at $i$-th step; $\psi_{i}$ is the $i$-th heading converted by quaternion iteration using equation (14) Diebel et al. (2006).
@@ 3.1 Dataset and Evaluation Metric
For the final evaluation of the proposed visual-inertial localization algorithm, we incorporated the Real-World Walk dataset, which includes two trajectories with time-aligned IMU data, discrete images, and ground truth positions. Additionally, 33 visual landmarks were employed as the VPR database.
Consistent with standard evaluation protocols for these datasets, we measured VPR model performance using Recall@N Arandjelovic et al. (2016). Additionally, the efficiency of the proposed VPR model was assessed through model parameter quantity and floating-point operations per second (FLOPS). GPS benchmarks were employed to evaluate the efficacy of our visual-inertial localization algorithm.
@@ 3.4 PDR-VPR Localization Performance Evaluation
  • Trajectory 2, Dense Feature Environment: This was conducted in a typical high-feature environment (see Figure 5), where the participant walked a closed loop among densely packed buildings. The environment offered numerous available visual features for positioning.
Figure 4: Trajectory comparison of Trajectory 1 in a sparse feature environment. The red points are correctly recognized, while the green ones are unrecognized. The indices of the green points from bottom to top are 1, 3, and 4. PDR-VPR w/ GES refers to the proposed visual-inertial localization algorithm; MDR-PDR represents the proposed dead-reckoning method.
Figure 5: Trajectory comparison of Trajectory 2 in dense feature environment. The red points are correctly recognized, while the green ones are unrecognized. The index of the green point is 1. PDR-VPR w/ GES refers to the proposed visual-inertial localization algorithm; MDR-PDR represents the proposed dead-reckoning method.
@@
We begin by presenting the evaluation results obtained in the lakeside environment, as illustrated in Figure 4. These results include both the proposed MDR-PDR trajectory and the fused PDR-VPR trajectory. In the sparse feature environment, we set a total of 11 visual beacons, of which 7 were successfully recognized (marked as Red Points) while 4 were unrecognized (marked as Green Points), as shown in Figure 6.
Figure 7: PDR-VPR without Gross Error Suppression for Trajectory 1
We aim to clarify the factors influencing this result. As shown in Figure 6, we provide a comparison of the actual landmark images corresponding to four unrecognized points. The unrecognized points may be attributed to several factors: the walking route actually passed through the visual beacons (Points 3 and 4), as well as a lack of distinctive features (Points 1 and 2). Additionally, to demonstrate the effectiveness and necessity of our gross error suppression method, we separately present the trajectory results corrected by the top-1 image from the proposed VPR model. As shown in Figure 7, without PDR prior information as a constraint, using image similarity alone as the recognition criterion leads to completely unusable localization results. This is due to the limited Recall@1 accuracy of the lightweight VPR network, particularly when the training and testing datasets are from completely different sources (our method was trained solely on the Pitts30k-train dataset). The results in Table 3 further support this phenomenon. Once a recognition error occurs, its impact on the subsequent localization results is catastrophic.
Figure 8: CDF Comparison for Trajectory 1. PDR-VPR w/ GES refers to the proposed visual-inertial localization algorithm; PDR-VPR w/o GES is the visual-inertial method without gross error suppression; MDR-PDR represents the proposed dead-reckoning method; PDR w/o MDR is the PDR algorithm without the MDR module.
With the aid of visual beacons, the fused trajectory aligns more closely with the ground truth compared to the MDR-PDR trajectory. This improvement is particularly evident in Figure 8, which displays the cumulative distribution function (CDF) curves for the fused trajectory (PDR-VPR w/ GES), fused trajectory without gross error suppression (PDR-VPR w/o GES), the MDR-PDR trajectory (MDR-PDR), and the PDR trajectory without MDR (PDR w/o MDR), highlighting the enhancements made.
Figure 9: CDF Comparison for Trajectory 2. PDR-VPR w/ GES refers to the proposed visual-inertial localization algorithm; PDR-VPR w/o GES is the visual-inertial method without gross error suppression; MDR-PDR represents the proposed dead-reckoning method; PDR w/o MDR is the PDR algorithm without the MDR module.
The proposed visual-inertial algorithm demonstrated improved performance in environments with dense features. As illustrated in Figure 5, out of 12 visual beacons, 11 were successfully recognized (indicated by the Red Points), while only one beacon failed to be recognized (shown as the Green Point), resulting in a recognition accuracy of 91.7%. The failed point is shown in the zoomed image in Figure 5, where the problematic beacon is situated among numerous repetitive trees, which obscure the nearby architectural features. Dense features are beneficial for more successful visual beacon recognition, and will contribute to enhancements in the fused trajectory. Figure 9 presents the CDF curves of the trajectories within the dense feature environment, showcasing these improvements.
@@ 4 Conclusion
This paper aims to provide stable and continuous positioning in large-scale urban environments by combining opportunistic visual landmark recognition and inertial pedestrian navigation, targeting potential applications on future mobile devices such as AR. To address this, we propose a lightweight MSGC-NetVLAD network for visual place recognition, which utilizes multi-scale convolutions and group convolutions to ensure a lightweight yet high-accuracy VPR. Additionally, a PDR method based on magnetic disturbance rejection is introduced to ensure positioning continuity. To address gross error issues in VPR, we propose using the natural divergence of the PDR system as a confidence threshold to constrain VPR observation updates. All modules are integrated using a Kalman filter framework. Extensive experiments demonstrate that our proposed MSGC-NetVLAD network achieves competitive accuracy. In comparison to the lightweight MobileNetV3-based VPR method, our proposed VPR solution enhances Recall@1 by at least 3% across two public datasets while reducing the number of parameters by 63.37%. Additionally, it achieves a performance level comparable to the larger VGG16-based method on the Pitts30k-test dataset, all with approximately 1.11 M parameters. Our overall PDR-VPR fusion algorithm improves the positioning accuracy of the proposed MDR-PDR by up to 46.86% at a confidence level of 75%.
However, the current work is limited to the algorithmic level, and we hope to integrate the proposed algorithm into mobile devices in the future. We believe this research will make a significant contribution to the application of AR technology and the development of smart cities.
@@ Acknowledgements
The authors would like to express their sincere gratitude to the editor and reviewers for their helpful comments throughout the review process, which significantly enhanced the quality of this paper. We also extend our thanks to Ms. Zhou Fengyi for her contributions to the design of the illustration graphs included in this work. Furthermore, I would like to thank Dr. Feng Chunying for conducting part of the experiments.